OpenStudy (anonymous):

Solve the inequality: 3x + 2 ≤ 1

OpenStudy (anonymous):

3x+2<=1 @mathmath333

OpenStudy (mathmath333):

we have done this before

OpenStudy (mathmath333):

just u have to shift 2 to RHS and then coefficient of x to RHS

OpenStudy (anonymous):

really

OpenStudy (mathmath333):

really

OpenStudy (anonymous):

this is different

OpenStudy (mathmath333):

give it a try

OpenStudy (anonymous):

ok so what do i do

OpenStudy (mathmath333):

shift 2 to RHS

OpenStudy (anonymous):

right hand side

OpenStudy (mathmath333):

yes

OpenStudy (anonymous):

3x<=1+2 @mathmath333

OpenStudy (mathmath333):

when u shift 2 to RHS also change the sign

OpenStudy (anonymous):

3x>=1+2

OpenStudy (mathmath333):

no i mean the sign in front of 2 not inequality sign

OpenStudy (anonymous):

3x<=1-2

OpenStudy (mathmath333):

\[\checkmark\]

OpenStudy (anonymous):

so what the inequality

OpenStudy (mathmath333):

what is \(1-2\)

OpenStudy (anonymous):

1

OpenStudy (anonymous):

X <1

OpenStudy (mathmath333):

1 is smaller than 2 so \(\Large 1-2=-1\)

OpenStudy (anonymous):

x<-1

OpenStudy (mathmath333):

\(\Large 3x<-1\)

OpenStudy (anonymous):

oh okay

OpenStudy (mathmath333):

but u have to put 3 now on RHS

OpenStudy (anonymous):

x<-1-3

OpenStudy (mathmath333):

no \(\Large 3\) is multilied with x u cannot do that

OpenStudy (anonymous):

ok so x<-1*3

OpenStudy (mathmath333):

no when a multiplication coefficent goes to RHS it becomes division

OpenStudy (anonymous):

x<-1/3

OpenStudy (mathmath333):

yes \(\Large x<-\dfrac{1}{3}\)

OpenStudy (mathmath333):

u need this http://www.mathsisfun.com/operation-order-pemdas.html